# Optimal Multi-Distribution Learning

### v2

@@ C.2.1 The existence of segments
For each $i\in\mathcal{W}^{j}$, there exists $1\leq s_{i}<e_{i}\leq T$, such that $\frac{1}{2^{j+2}}<w^{s_{i}}_{i}\leq\frac{1}{2^{j+1}}$, $\frac{1}{2^{j}}<w^{e_{i}}_{i}$ and $w^{t}_{i}>2^{-(j+2)}$ for any $s_{i}\leq t\leq e_{i}$. In words, there exists a $\left(\frac{1}{2^{j+1}},\frac{1}{2^{j+2}},\log(2)\right)$-segment $(s_{i},e_{i})$ with the index set as $\{i\}$.
@@
By definition, it is easy to find $e_{i}$. To find $s_{i}$, define $\tau=\max\{t\leq e_{i}|w^{t}_{i}\leq 2^{-(j+2)}\}$. $\tau$ is properly defined because $w^{1}_{i}=\frac{1}{k}\leq 2^{-(j+2)}$. By definition, we have that for any $\tau+1\leq t\leq e_{i}$, $w^{t}_{i}>2^{-(j+2)}$. Noting that $\log_{2}(w^{t+1}_{i}/w^{t}_{i})\leq\frac{\eta-\log_{2}(Z_{t+1}/Z_{t})}{\log(2)}\leq\frac{2\eta}{\log(2)}\leq\frac{1}{4}$, we have that $w^{\tau+1}_{i}\leq\frac{1}{2^{j+2}}\cdot 2^{\frac{1}{4}}\leq\frac{1}{2^{j+1}}$. So it suffices to choose $s_{i}=\tau+1$. ∎
C.2.2 Extracting regular segments from irregular segments
@@
Recall the definition of $\mathcal{W}_{j}$ and $(s_{i},e_{i})$ for $i\in\mathcal{W}_{j}$ in Lemma 26. Then there exists a group of subsets $\{\mathcal{W}_{j}^{p}\}_{p=1}^{P}$ of $\mathcal{W}_{j}$ such that
  1. (i).
    $\mathcal{W}_{j}^{p}\subset\mathcal{W}_{j}$, $\mathcal{W}_{j}^{p}\cap\mathcal{W}_{j}^{p^{\prime}}=\emptyset$, $\forall p\neq p^{\prime}$;
  2. (ii).
  3. (iii).
    There exists $1\leq\tilde{s}_{1}<\tilde{e}_{1}\leq\tilde{s}_{2}<\tilde{e}_{2}\leq\dots\leq\tilde{s}_{P}<\tilde{e}_{P}\leq T$, such that for each $1\leq p\leq P$, $\tilde{s}_{p}=\min_{i\in\mathcal{W}_{j}^{p}}s_{i},\tilde{e}_{p}=\max_{i\in\mathcal{W}_{j}^{p}}e_{i}$ and $\max_{i\in\mathcal{W}_{j}^{p}}s_{i}\leq\min_{i\in\mathcal{W}_{j}^{p}}e_{i}$.

### v3

@@ D.1 Proof of Theorem 2
Consider $\widetilde{i}\in[k]$. If $\mathbb{P}_{\mathcal{G}^{\prime}}\{h_{\mathrm{out}}=h^{\star},M_{i}\leq m\}\geq 1/2$ for some $m\geq 0$, then for any $h\in\mathcal{H}_{\widetilde{i}}$, one has
$\mathbb{P}_{\mathcal{G}^{\prime}}\big{\{}h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m\big{\}}\geq\frac{1}{2}\mathbb{P}_{\mathcal{G}^{\prime}}\big{\{}h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\big{\}}\exp\big{(}-80\sqrt{m}\varepsilon-40m\varepsilon^{2}\big{)},$
and moreover, $m$ necessarily exceeds $m\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}$.
@@
See Appendix D.2. ∎
In view of Lemma 18, we have
$\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}<\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right\}<\frac{1}{2}.$ (124)
Observing that $\mathbb{P}_{\mathcal{G}^{\prime}}\{h_{\mathrm{out}}=h^{\star}\}\geq 3/4$, we can derive
$\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right\}$ $=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star}\right\}-\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}<\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right\}$
$>\frac{3}{4}-\frac{1}{2}=\frac{1}{4},$
which implies that
$\mathbb{E}_{\mathcal{G}^{\prime}}[M_{\widetilde{i}}]$ $\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\cdot\mathbb{P}_{\mathcal{G}^{\prime}}\left\{M_{\widetilde{i}}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right\}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\cdot\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}\right\}$
$\geq\frac{\log(N_{0}/4)}{120000\varepsilon^{2}}\geq\frac{d-\log_{2}(8k)}{120000\varepsilon^{2}}\geq\frac{d}{240000\varepsilon^{2}}.$
Summing over $i\in[k]$ gives
$\mathbb{E}_{\mathcal{G}^{\prime}}\left[\sum_{i=1}^{k}M_{\widetilde{i}}\right]\geq\frac{dk}{240000\varepsilon^{2}},$ (125)
thereby concluding the proof.
D.2 Proof of Lemma 18
Consider any $h\in\mathcal{H}_{\widetilde{i}}$ (and hence $h\neq h^{\star}$). For $v=[v_{p}]_{1\leq p\leq m}\in\{-1,1\}^{m}$, we let
$n^{+}(v)=\sum_{p=1}^{m}\mathds{1}\{v_{p}=1\}$
denote the number of $1$’s in the coordinates of $v$. Let $\mathcal{V}$ be a subset of $\{-1,1\}^{2m}$ defined as
$\mathcal{V}\coloneqq\big{\{}v^{1},v^{2}\in\{-1,1\}^{m}\mid n^{+}(v^{1})-n^{+}(v^{2})\leq 4\sqrt{m}+2m\varepsilon\big{\}}.$
Let $\mathbb{P}_{\mathcal{C}}\{\cdot\}$ denote the probability distribution of $\left\{\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}\right\}$, and $\mathbb{P}_{\mathcal{C}^{\prime}}\{\cdot\}$ the probability distribution of $\left\{\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}\right\}$. Hoeffding’s inequality then tells us that
$\mathbb{P}_{\mathcal{C}}\{\mathcal{V}\}\geq\frac{3}{4}.$
Also, observe that
$\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m,\left\{\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}\right\}\in\mathcal{V}\right\}$
$\qquad\geq\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\right\}-\left(1-\mathbb{P}_{\mathcal{C}}\left\{\mathcal{V}\right\}\right)\geq\frac{3}{4}-\left(1-\frac{3}{4}\right)=\frac{1}{2},$ (126)
namely,
$\sum_{v=\{v^{1},v^{2}\}\in\mathcal{V}}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m,\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}\geq\frac{1}{2}.$
In addition, for any $v=\{v^{1},v^{2}\}\in\mathcal{V}$, it is readily seen that
$\mathbb{P}_{\mathcal{C}^{\prime}}\{v\}$ $=\mathbb{P}_{\mathcal{C}}\{v\}\cdot\left(1-8\varepsilon\right)^{n^{+}(v^{1})-n^{+}(v^{2})}(1+8\varepsilon)^{n^{+}(v^{2})-n^{+}(v^{1})}$
$=\mathbb{P}_{\mathcal{C}}\{v\}\left(\frac{1-8\varepsilon}{1+8\varepsilon}\right)^{n^{+}(v^{1})-n^{+}(v^{2})}$
$\geq\mathbb{P}_{\mathcal{C}}\{v\}\exp\Big{(}-20\big{(}n^{+}(v^{1})-n^{+}(v^{2})\big{)}\varepsilon\Big{)}$
$\geq\mathbb{P}_{\mathcal{C}}\{v\}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2}),$
where the last line follows from the definition of $\mathcal{V}$. As a result, we can demonstrate that, for any $v=\{v^{1},v^{2}\}\in\mathcal{V}$,
$\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m,\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$
$=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m\mid\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}\mathbb{P}_{\mathcal{C^{\prime}}}\{v\}$
$\geq\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m\mid\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}\mathbb{P}_{\mathcal{C}}\{v\}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2})$
$=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\mid\{x_{j_{h,\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{2}\right\}\mathbb{P}_{\mathcal{C}}\{v\}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2})$ (127)
$=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m,\{x_{j_{h,\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{2}\right\}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2}).$
To see why (127) holds, observe that (here, for any $1\leq l\leq m$, let $v^{1}_{l}$ (resp. $v^{2}_{l}$) be the $l$-th coordinate of $v^{1}$ (resp. $v^{2}$)):
$\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m\mid\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$
$=\sum_{m^{\prime}=1}^{m}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}=m^{\prime}\mid\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$
$=\sum_{m^{\prime}=1}^{m}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}=m^{\prime}\mid\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m^{\prime}}=\{v^{1}_{l}\}_{l=1}^{m^{\prime}},\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m^{\prime}}=\{v^{2}_{l}\}_{l=1}^{m^{\prime}}\right\}$ (128)
$=\sum_{m^{\prime}=1}^{m}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}=m^{\prime}\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m^{\prime}}=\{v^{1}_{l}\}_{l=1}^{m^{\prime}},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m^{\prime}}=\{v^{2}_{l}\}_{l=1}^{m^{\prime}}\right\}$ (129)
$=\sum_{m^{\prime}=1}^{m}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}=m^{\prime}\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$ (130)
$=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}.$
where (129) results from Lemma 19 in Appendix D.3, and (128) and (130) hold since for $h^{\prime}=h,h^{\star}$, the event $\{h_{\mathrm{out}}=h^{\prime},M_{\widetilde{i}}=m^{\prime}\}$ is independent of $\{x^{l}(\widetilde{i})\}_{l\geq m^{\prime}+1}$. Taking the sum over $v=\{v^{1},v^{2}\}\in\mathcal{V}$, we obtain
$\mathbb{P}_{\mathcal{G}^{\prime}}\{h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m\}$
$\geq\sum_{v\in\mathcal{V}}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{\widetilde{i}}\leq m,\{x_{j_{h^{\star},\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h,\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{2}\right\}$
$\geq\sum_{v\in\mathcal{V}}\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m,\{x_{j_{h,\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}=v^{2}\right\}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2})$
$=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m,\left\{\{x_{j_{h,\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m},\{x_{j_{h^{\star},\widetilde{i}}}^{\ell}(\widetilde{i})\}_{\ell=1}^{m}\right\}\in\mathcal{V}\right\}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2})$
$\geq\frac{1}{2}\mathbb{P}_{\mathcal{G}^{\prime}}\big{\{}h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\big{\}}\exp(-80\sqrt{m}\varepsilon-40m\varepsilon^{2}),$
where the last line arises from (126).
Summing over all $h\in\mathcal{H}_{\widetilde{i}}$, we reach
$1\geq\mathbb{P}_{\mathcal{G}^{\prime}}\big{\{}h_{\mathrm{out}}\in\mathcal{H}_{\widetilde{i}},M_{\widetilde{i}}\leq m\big{\}}\geq\frac{N_{0}}{2}\mathbb{P}_{\mathcal{G}^{\prime}}\big{\{}h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\big{\}}\exp(-40m\varepsilon^{2}-80\sqrt{m}\varepsilon),$ (131)
given that each $\mathcal{H}_{\widetilde{i}}$ contains $N_{0}$ hypotheses. This in turn reveals that
$\frac{1}{2}\leq\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{\widetilde{i}}\leq m\right\}\leq\frac{2}{N_{0}}\exp(40m\varepsilon^{2}+80\sqrt{m}\varepsilon).$ (132)
Consequently, we arrive at
$40m\varepsilon^{2}+80\sqrt{m}\varepsilon\geq\log(N_{0}/4),$
which implies that
$m\geq\min\bigg{\{}\frac{\log(N_{0}/4)}{80\varepsilon^{2}},\frac{\log^{2}(N_{0}/4)}{30000\varepsilon^{2}}\bigg{\}}\geq\frac{\log(N_{0}/4)}{30000\varepsilon^{2}}.$
D.3 Statement and proof of Lemma 19
@@
For any $i\in[k]$ and $l\geq 1$, let $x^{l}(i)$ denote the $l$-th sample from $\mathcal{D}_{i}$. For any $\widetilde{i}\in[k]$, $h\in\mathcal{H}_{\widetilde{i}}$, $m>0$, and $v^{1},v^{2}\in\{-1,1\}^{2m}$, one has
$\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h^{\star},M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2}\right\}$
$=\mathbb{P}_{\mathcal{G}^{\prime}}\left\{h_{\mathrm{out}}=h,M_{i}=m\mid\{x_{j_{h,\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{2},\{x_{j_{h^{\star},\widetilde{i}}}^{l}(\widetilde{i})\}_{l=1}^{m}=v^{1}\right\}.$